# An Ran Chen

### v1

@@ 5 Case Study Results
in which m represents each of the top 5 methods in the Stack Trace, and $CL_{m}$ is the number of lines in the correspondent method that is covered by the test. Then, we select the set of the X highest results by applying the following formula, in which T represents the set of tests and X is a threshold for the number of selected tests:
$\begin{split}T_{\text{failing}}=\{t\in T\mid t\text{ is one of the }X\text{ tests with the highest }\\ ST\_Covered\_lines\_number_{t}\}\end{split}$ (3)
We noticed that selecting a value for X that was too low made the approach often miss buggy locations because it was too focused on just a few methods. On the other hand, a high threshold also generates inaccurate results since the approach is unable to differentiate between the methods in the stack trace. Upon experimentation, we noticed that setting it to 15 gave the best results. Finally, we apply the Ochiai formula (1) using these 15 tests as failing tests and all the remaining as passing tests.
Stack Trace (ST) Score. The SB score described above merges the information in the stack traces with the test coverage to compute the suspiciousness score for a given method. In addition to that, we make use of the ST score used in a previous study (Chen et al., 2021b) in order to boost the stack trace’s impact into the calculated score. To do this, after having computed the SB_score following the approach described before, we sum to it a ST_score calculated in the following manner:
@@
In which ST_rank is the position in which the given method appears in the stack trace after the external entries are removed. For example, if the method appears in the second position, the ST_score would be 1/2=0.5.
Suspiciouness Score. After having both the STCB Ochiai score and the ST Score, we compute the suspiciousness score of a giving method by applying the formula:
$Suspiciouness\_score=STCB\_Ochiai\_score+ST\_score$ (5)
Suspiciouness Rank. After having the final suspiciousness score, we generate the final suspiciousness rankings. The higher the suspiciousness score, the higher the ranking. For instance, if the highest calculated suspiciousness score is 1.0 for the method M8, M8 is going to be the first method in the suspiciousness rank, as it has the highest probability of containing the bug.
@@
      Closure (8)       Ochiai       0       0       0       0.0007       0.0012
      Stack Trace       1       2       2       0.1710       0.2170
@@
      Stack Trace       1       1       1       0.3333       1
      SB_score       1       1       1       0.3893       1
      SBEST       1       1       1       0.5853       1
      Compress (11)       Ochiai       0       0       1       0.0079       0.023
      Stack Trace       3       6       8       0.4122       0.4425
      SB_score       0       3       3       0.1404       0.1587
      SBEST       3       9       9       0.4620       0.4904
@@
      Stack Trace       1       2       2       0.5196       0.5196
      SB_score       1       1       1       0.3338       0.3338
      SBEST       1       2       2       0.5101       0.5101
      JacksonCore (4)       Ochiai       0       1       1       0.0693       0.0836
@@
      SB_score       0       0       0       0.0409       0.0457
      SBEST       0       1       1       0.1609       0.1634
      JacksonDatabind (10)       Ochiai       0       0       0       0.0005       0.0005
      Stack Trace       1       2       3       0.1686       0.2446
      SB_score       0       1       1       0.0125       0.0212
      SBEST       1       3       3       0.0359       0.0777
@@
      Stack Trace       3       3       5       0.4410       0.4444
      SB_score       0       0       0       0.0214       0.0318
      SBEST       3       4       4       0.4070       0.4279
@@
      Stack Trace       1       1       1       0.3333       0.3333
      SB_score       0       0       0       0.0625       0.0625
@@
      Lang (4)       Ochiai       0       0       0       0.0008       0.0008
      Stack Trace       1       3       3       0.4375       0.4375
      SB_score       0       1       1       0.1411       0.1411
      SBEST       2       3       3       0.5835       0.5835
@@
      Stack Trace       2       2       2       0.2215       0.4583
      SB_score       0       0       0       0.0428       0.0292
      SBEST       2       2       2       0.1372       0.3472
      Mockito (2)       Ochiai       0       0       0       0.0060       0.0281
      Stack Trace       0       1       1       0.0238       0.1667
      SB_score       0       1       1       0.0371       0.25
      SBEST       0       1       1       0.0371       0.25
@@
      Stack Trace       0       1       2       0.2917       0.2917
      SB_score       1       1       1       0.5001       0.5001
      SBEST       0       1       1       0.3333       0.3333
      Total (60)       Ochiai       0       1       2       0.00902       0.01421
      Stack Trace       16       27       34       0.32406       0.40992
      SB_score       4       10       11       0.20396       0.26517
      SBEST       17       32       33       0.42846       0.49647
@@
Results. The Stack Traces ranking alone locates 34 out of the 60 bugs in the Top-5. Table 5 presents the metrics results for the 2 baselines as well as for the SB_score alone and the SBEST approach for all the systems. We can see that the worst-performing technique is Ochiai, which, as discussed before, locates only 2 bugs in Top-5. The best-performing approach is SBEST, but the Stack Traces ranking is a close second. The Stack Trace ranking locates 16 bugs in Top-1, 27 in Top-3 and 34 in Top-5. Surprisingly, the Stack Traces alone provide very good results, locating more than half of the bugs (56.67%) within the Top 5. This shows that the stack traces are a very important source of information about the bug location and should be prioritized when it comes to studying FL in crash reports. The SB_score is able to locate some bugs that the Stack Trace ranking does not locate, but in general, it does not perform very well, identifying only 18.33% of the buggy methods in the Top 5. We can see, however, that the SB_score is substantially improved by the addition of the ST_score. The resultant technique SBEST performs slightly better than the Stack Traces in the Top 1 and Top 3, locating 17 and 32 bugs, respectively, and locates only one bug less in the Top 5, totalling 33 bugs spotted within the Top 5 rank. This technique and the Stack Traces ranking take turns as the best localization approach for each project. For example, the Stack Trace ranking is better overall for the JacksonCore project, while SBEST has the top performance on the Lang bugs. In total, SBEST locates 17 bugs in Top-1, 32 in Top-3 and 33 in Top-5. The average MAP and MRR for SBEST are 0.42846 and 0.49647, respectively, representing a 32.22% improvement on the MAP and a 17.43% in the MRR when compared to the Stack Traces. These results show that the stack traces can complement existing passing test information to help locate the bug root causes.
@@
Overall, out of 60 studied bugs, SBEST successfully located 17 bugs in Top-1, 32 in Top-3 and 33 in Top-5. In addition, it achieves an improvement of 32.22% on MAP and 17.43% on MRR when compared to the Stack Trace ranking.
6 Discussion and Implication of Our Findings
@@
On the use of stack traces in the absence of failing tests. From the results obtained in Section 5, we observe that the majority of the crash report bugs do not contain fault-triggering tests. In such cases, the test failure information is not available for the fault localization process, causing techniques such as SBFL to fail. The stack traces, on the other hand, demonstrate to be deeply correlated to the bug cause. We found that, in 66% of the bugs, at least one buggy method is directly listed in the stack trace. Even when this is not the case, the buggy methods are usually just a few calls away from it. In addition, the most common bugfix intention is to prevent the exception in the stack trace from happening again, which highlights the big association between them.
@@
Leveraging stack trace rankings for enhanced fault localization. In the real world, test failures are not always present. Especially when it comes to production-phase bugs, it is important to look for alternative sources of information for fault localization. Making use of the data available in logs and stack traces shows to be a good path to take in such scenarios. Even though previous studies have used stack traces for Fault Localization, many are based on information retrieval (IRFL) approaches (Zhou et al., 2012; Saha et al., 2013; Lam et al., 2017). IRFL techniques treat all the content in the bug report, including the stack traces, as textual information, therefore missing important context such as the ranking of the stack traces. Our results show that the stack traces ranking alone was able to locate more than half of the bugs within Top-5 and, thus, should be better utilized.
@@
Encouraging the integration of stack traces to complement execution information. SBEST, our approach combining the stack trace ranking with the code coverage, shows promising results. The improvement in the metrics implies that the coverage information was able to enhance the stack trace ranking, aiding in obtaining a better overview of the system execution in the moment of failure. Our main goal when studying this technique was to gain a deeper understanding of the scenario of these bugs and how each piece of information available can help with fault localization. We believe that this is just the first step and that future research can benefit from the aforementioned findings to build more intricate techniques. In addition, we make all the data from this study available at Zenodo [^1] https://zenodo.org/records/11062413, including the detailed code coverage and test results of the bug report commits that took us 100 hours to obtain. We think that this data will be useful for future research on subjects such as fault localization, automatic bug fixing, test generation, etc. Our data does not contain post-bug-fixing development knowledge, which will allow researchers to tailor tools more aligned with the real-world scenario.
@@ 7 Threats to Validity
Threats to external validity relate to how generalizable our findings are. To minimize this threat, we conduct our studies in 15 systems from Defects4j, a famous framework of real bugs used in multiple other studies. These systems are widely used and vary in size and number of tests. Even though every system is written in Java, our method is not restricted to Java-based systems. In addition, despite the low percentage of bugs with stack traces in Defects4j, this aligns with the percentage reported in other systems (Chen et al., 2021a). Moreover, we opted not to use the buggy version made available in defects4j but the code available at the moment of the bug report creation in order to better emulate the scenario in which the developers investigate the bug. Finally, even though we opted for applying only Ochiai when analyzing SBFL’s efficiency when applied to crash report bugs, it is shown to be one of the best SBFL techniques and performs very well on real faults (Pearson et al., 2017).
@@
Threats to internal validity relate to the extent to which the results of the study can be attributed solely to the experimental treatments and not to flaws in the experimental design. In this study, we only analyze Defects4j bugs. Although Defects4j is widely used (Lutellier et al., 2020; Chen et al., 2021b; Li et al., 2021; Ye et al., 2022), the findings might be different in other systems, specially non-Java systems. Another threat to internal validity is the way in which we selected the bug reports with stack traces. We collect stack traces in the description and comments, which is the typical place in which they are added. However, in some cases, developers may also add stack traces via attachments. Nonetheless, this is very rare, representing less than 1% of the cases based on a previous study from Chen et al. (2021a).
@@ 8 Conclusion
Spectrum-based fault Localization (SBFL) approaches are highly recognized due to their accuracy, efficiency, and simplicity. However, their effectiveness depends on the availability of fault-triggering failing tests. In the lack of such test failures, alternative approaches should be considered. In this paper, we research using stack traces for SBFL in the absence of failing tests. We study 60 crash reports from the Defects4J benchmark. We find that most of the studied crash reports do not contain fault-triggering failing tests. This results in very low efficiency when using traditional SBFL. On the other hand, when it comes to the stack traces, we find that in 98.3% of the studied bugs, the bugfix intention is directly correlated with the exception in the stack trace. In addition, in 78.3% of the bugs, the buggy method is directly reachable from the stack trace, with an average distance of only 0.34 method calls. This shows that the buggy methods are usually very close to the stack traces when it comes to the execution call graph. On top of that, our results show that even without any advanced technique, the stack traces alone provide a good indication of the buggy locations, being capable of locating more than half of the bugs on Top-5. Finally, we develop a simplified SBFL method called SBEST that uses the stack trace information in place of the failing tests and thus integrates it into the coverage information to perform the SBFL. SBEST was able to locate 17 bugs in Top-1, 32 in Top-3 and 33 in Top-5, which represents an improvement of 32.22% on MAP and 17.43% on MRR when compared to the Stack Trace ranking. The enhancement of the metrics suggests that the coverage information improved the stack trace ranking and helped the technique gain a clearer picture of the system’s execution at the time of failure. We believe that future studies can benefit from the findings of this research to develop more intricate techniques. In addition, all the data that we have made available can help studies in areas such as automatic bug fixing, test generation, and, of course, fault localization. As potential future work, we plan to expand our analysis by increasing our bugs dataset utilizing other sources beyond Defects4j.

### v2

@@ 4 Results and Analysis
in which m represents each of the top 5 methods in the Stack Trace, and $CL_{m}$ is the number of lines in the corresponding method covered by the test. Then, we select the set of X highest results by applying the following formula, in which T represents the set of tests and X is a threshold for the number of selected tests:
[…]
We noticed that selecting a value for X that was too low often missed buggy locations because it was too focused on just a few methods. On the other hand, a high threshold also generates inaccurate results since the approach cannot differentiate between the methods in the stack trace. Experimenting showed that setting it to 15 gave the best results. Finally, we apply the Ochiai formula (1) using these 15 tests as failing tests and all the remaining as passing tests. Table 6 presents the experimental results for different values of proxy failing tests X, which will be discussed further in the results section.
Stack Trace (ST) Score. The SB score described above merges the information in the stack traces with the test coverage to compute the suspiciousness score for a given method. In addition, we use the ST score used in a previous study (Chen et al., 2021b) to boost the stack trace’s impact on the calculated score. To do this, after having computed the SB_score following the approach described before, we add to it a ST_score calculated in the following manner:
@@
In which ST_rank is the position where the given method appears in the stack trace after the external entries are removed. For example, if the method appears in the second position, the ST_score would be 1/2=0.5.
Suspiciousness Score. After having both the SB Ochiai score and the ST Score, we compute the suspiciousness score of a given method by applying the formula:
Suspiciousness_score=SB_Ochiai_score+ST_scoreSuspiciousness\_score=SB\_Ochiai\_score+ST\_score (5)
Suspiciousness Rank. After having the final suspiciousness score, we generate the final suspiciousness rankings. The higher the suspiciousness score, the higher the ranking. For instance, if the highest calculated suspiciousness score is 1.0 for the method M8, M8 will be the first method in the suspiciousness rank, as it has the highest probability of containing the bug.
@@
   Closure (8)    Ochiai    0    0    0    0.001    0.001
   Stack Trace    1    2    2    0.171    0.217
@@
   Stack Trace    1    1    1    0.333    1
   SB_score    1    1    1    0.389    1
   CFBFL    1    1    1    0.399    1
   SBEST    1    1    1    0.585    1
   Compress (11)    Ochiai    0    0    1    0.008    0.023
   Stack Trace    3    6    8    0.412    0.443
   SB_score    0    3    3    0.141    0.159
   CFBFL    0    1    1    0.002    0.002
   SBEST    3    9    9    0.462    0.491
@@
   Stack Trace    1    2    2    0.52    0.52
   SB_score    1    1    1    0.334    0.334
   CFBFL    0    0    0    0.0001    0.0001
   SBEST    1    2    2    0.51    0.51
   JacksonCore (4)    Ochiai    0    1    1    0.069    0.084
@@
   SB_score    0    0    0    0.041    0.046
   CFBFL    0    0    1    0.001    0.001
   SBEST    0    1    1    0.161    0.164
   JacksonDatabind (10)    Ochiai    0    0    0    0.001    0.001
   Stack Trace    1    2    3    0.169    0.245
   SB_score    0    1    1    0.013    0.021
   CFBFL    1    1    1    0.176    0.215
   SBEST    1    3    3    0.036    0.078
@@
   Stack Trace    3    3    5    0.441    0.445
   SB_score    0    0    0    0.022    0.032
   CFBFL    2    3    4    0.224    0.299
   SBEST    3    4    4    0.407    0.428
@@
   Stack Trace    1    1    1    0.333    0.333
   SB_score    0    0    0    0.063    0.063
   CFBFL    0    0    0    0.001    0.001
@@
   Lang (4)    Ochiai    0    0    0    0.001    0.001
   Stack Trace    1    3    3    0.438    0.438
   SB_score    0    1    1    0.141    0.141
   CFBFL    0    0    1    0.01    0.01
   SBEST    2    3    3    0.584    0.584
@@
   Stack Trace    2    2    2    0.222    0.458
   SB_score    0    0    0    0.041    0.029
   CFBFL    0    1    1    0.113    0.129
   SBEST    2    2    2    0.137    0.347
   Mockito (2)    Ochiai    0    0    0    0.006    0.028
   Stack Trace    0    1    1    0.024    0.167
   SB_score    0    1    1    0.037    0.25
   CFBFL    0    0    0    0.002    0.002
   SBEST    0    1    1    0.037    0.25
@@
   Stack Trace    0    1    2    0.292    0.292
   SB_score    1    1    1    0.5    0.5
   CFBFL    0    0    0    0.001    0.001
   SBEST    0    1    1    0.333    0.333
   Total (60)    Ochiai    0    1    2    0.009    0.014
   Stack Trace    16    27    34    0.324    0.41
   SB_score    4    10    11    0.204    0.265
   CFBFL    4    8    13    0.084    0.136
   SBEST    17    32    33    0.429    0.497
Table 6: Fault Localization Performance with Different Numbers of Failing Tests for SBEST
   X (Proxy Failing Tests)    Top-1    Top-3    Top-5    MAP    MRR
   5    14    22    27    0.37111    0.43887
   10    16    29    30    0.39121    0.43944
   15    17    32    33    0.42846    0.49647
   20    14    23    28    0.37532    0.42598
   25    12    20    25    0.30897    0.41720
Table 7: Fault Localization Performance for Different Numbers of Top Stack-Trace Methods for SBEST
Top #Stack-trace Methods Top-1 Top-3 Top-5 MAP MRR
5 17 32 33 0.42846 0.49647
10 7 10 11 0.38962 0.41266
15 4 7 9 0.20011 0.2409
@@
Results. The Stack Traces ranking alone locates 34 out of the 60 bugs in the Top-5. Table 5 presents the metrics results for the 2 baselines as well as for the SB_score alone and the SBEST approach for all the systems. We can see that the worst-performing technique is Ochiai, which, as discussed before, locates only 2 bugs in the Top 5. The best-performing approach is SBEST, but the Stack Traces ranking is a close second. The Stack Trace ranking locates 16 bugs in Top-1, 27 in Top-3, and 34 in Top-5. Surprisingly, the Stack Traces alone provide very good results, locating more than half of the bugs (56.67%) within the Top 5. This shows that stack traces are a significant source of information about bug location and should be prioritized when studying FL in crash reports. The SB_score can locate some bugs that the Stack Trace ranking does not locate, but in general, it does not perform very well, identifying only 18.33% of the buggy methods in the Top 5.
CFBFL, which ranks methods based on their frequency in unique stack contexts during execution, identified 4 bugs in Top-1, 8 in Top-3, and 13 in Top-5. While it improves over Ochiai, it lags behind Stack Trace ranking and SBEST in all metrics. The MAP and MRR of CFBFL (0.084 and 0.136) are significantly lower than those of SBEST and Stack Traces, suggesting that while call frequency provides useful contextual signals, it may not be sufficient on its own for localizing crash faults when fault-triggering test coverage is sparse or missing.
We observe that adding the ST score improves the SB_score. The resulting technique, SBEST, outperforms the Stack Traces approach in the Top 1 and Top 3 ranks, locating 17 and 32 bugs (compared to Stack Traces’ 16 and 27), and only one fewer bug in the Top 5, for a total of 33. Overall, SBEST and Stack Traces are the two most effective localization techniques across all projects. Stack Traces performs better on JacksonCore, while SBEST achieves the best results on Lang. On average, SBEST attains a MAP of 0.42846 and an MRR of 0.49647, showing moderate improvements (32.22% and 17.43%) over the Stack Trace baseline in scenarios where fault-triggering tests are unavailable. These results indicate that stack traces can effectively complement passing test information in identifying root causes.
@@
Overall, out of the 60 studied bugs, SBEST successfully located 17 bugs in the Top 1, 32 in the Top 3, and 33 in the Top 5. Additionally, it achieves improvements of 32.22% on MAP and 17.43% on MRR compared to the Stack-Trace-based ranking.
SBEST performs best with 15 proxy failing tests and the top 5 stack trace methods. Tables 6 and 7 present the fault localization performance of SBEST under different parameter settings for the number of proxy failing tests and the number of stack trace methods, respectively. FL performance is stable between 10-20 proxy failing tests, with Top-1 results peaking at 17 faults. However, the performance degrades when we include more than 15 proxy failing tests or more than 5 stack trace methods. This suggests that, when too many failing tests or stack-trace methods are included, more noise may be introduced, which can negatively impact fault localization accuracy. For example, using 25 proxy failing tests or 15 stack methods significantly reduces Top-1 and MAP scores.
Setting the number of proxy failing tests (X) to 15 and using the top 5 stack trace methods gives the best results, with 17 bugs in Top-1 and a MAP of 0.4285. Other configurations show a drop in performance.
5 Discussion and Implications of Our Findings
@@
On the use of stack traces in the absence of failing tests. From the results obtained in Section 4, we observe that most crash report bugs do not contain fault-triggering tests. In such cases, the test failure information is unavailable for the fault localization process, causing techniques such as SBFL to fail. Conversely, the stack traces are deeply correlated with the cause of the bug. We found that, in 66% of the bugs, at least one buggy method is directly listed in the stack trace. Even when this is not the case, the buggy methods are usually just a few calls away from it. In addition, the most common bugfix intention is to prevent the exception in the stack trace from happening again, highlighting the significant association between them.
@@
Leveraging stack trace rankings for enhanced fault localization. In the real world, test failures are not always present. Especially regarding production-phase bugs, it is essential to look for alternative sources of information for fault localization. Using the data available in logs and stack traces is a good path for such scenarios. Although previous studies have utilized stack traces for Fault Localization, many are based on information retrieval (IRFL) approaches (Zhou et al., 2012; Saha et al., 2013; Lam et al., 2017). IRFL techniques treat all content in the bug report, including stack traces, as textual information; therefore, they miss important context, such as the ranking of the stack traces. Our results show that the stack traces ranking alone could locate more than half of the bugs within Top-5 and, thus, should be better utilized.
@@
Encouraging the use of stack traces to complement execution information. SBEST, our approach combining the stack trace ranking with the code coverage, shows promising results. The improvement in the metrics implies that the coverage information enhanced the stack trace ranking, aiding in obtaining a better overview of the system execution at the moment of failure. Our primary goal in studying this technique was to gain a deeper understanding of the scenario surrounding these bugs and how each piece of available information can aid in fault localization. We believe this is just the first step, and that future research can benefit from the findings above to develop more sophisticated techniques. In addition, we make all the data from this study available at Zenodo [^1] https://zenodo.org/records/11062413, including the detailed code coverage and test results of the bug report commits that took us 100 hours to obtain. We believe this data will be beneficial for future research on fault localization, automatic bug fixing, test generation, and other related areas. Our data excludes any post-fix developer knowledge, which would allow researchers to tailor tools more closely aligned with real-world scenarios.
Limitations. While our findings highlight the potential of stack traces for fault localization in the absence of fault-triggering tests, we acknowledge several limitations. Our study focuses specifically on bugs that result in exceptions with available stack traces. In practice, many bugs do not manifest as crashes or produce exceptions, which may limit the applicability of our findings to the broader space of software faults. Future work should investigate the applicability of stack-trace-based localization to a wider range of bug types and software systems.
Future Directions. Our findings open several avenues for future research. One particularly important direction is to explore how execution signals such as stack traces and runtime logs can be better leveraged in the absence of fault-triggering tests. These lightweight signals can serve as a proxy for fault-triggering execution, allowing fault localization methods like SBFL to operate in real-world scenarios where fault-triggering tests are missing. Additionally, future studies could expand evaluation beyond Defects4J to improve generalizability across projects and programming languages. Some recent research, such as AutoFL (Kang et al., 2024b), applies LLMs to reason over code and coverage data for fault localization. At the same time, LIBRO (Kang et al., 2024a) generates bug-reproducing tests directly from natural language bug reports with stack traces. Building on these ideas, one direction is to leverage LLMs to interpret the underlying causes of failures from stack traces and coverage data, and use that information to filter out irrelevant methods or code regions. This could enable more accurate and efficient ranking of suspicious methods, potentially supporting automated patch generation and thereby enhancing the practicality of SBEST in scenarios where explicit test failures are absent.
@@ 6 Threats to Validity
Threats to external validity relate to the generalizability of our findings. To minimize this threat, we selected 15 systems from Defects4J, as used by  Just et al. (2014b), a widely used benchmark of real bugs in prior studies. These systems vary in size, test coverage, and domain. However, we acknowledge that our study focuses on 60 crash-report bugs ($\approx$10% of Defects4J) with available stack traces. This proportion aligns with observations from real-world studies, which show that only a limited subset of user-reported bugs contain stack traces (Bettenburg et al., 2008; Chen et al., 2021a). Although this represents a subset of all bugs, crash reports are among the most practically significant cases developers face, as stack traces often provide the first diagnostic evidence when failures occur and have been widely utilized in fault localization research (Bettenburg et al., 2008; Zou et al., 2021). We prioritized crash-reporting bugs for which stack traces and test coverage could be reliably extracted, though this was sometimes challenging due to legacy and dependency issues. We publicly release our dataset to support replication and further study. Finally, while our evaluation focuses on Java projects, the proposed approach can be extended to other programming languages in future work.
@@
Threats to internal validity relate to the extent to which the results of the study can be attributed solely to the experimental treatments and not to flaws in the experimental design. In this study, we only analyze Defects4J bugs. Although Defects4J is widely used (Lutellier et al., 2020; Chen et al., 2021b; Li et al., 2021; Ye et al., 2022), the findings might be different in other systems, especially non-Java systems. Another threat to internal validity is how we selected the bug reports with stack traces. We collect stack traces in the description and comments, which is the typical place in which they are added. However, in some cases, developers may also add stack traces via attachments. Nonetheless, this is very rare, representing less than 1% of the cases based on a previous study from Chen et al. (2021a).
@@ 7 Conclusion
Spectrum-based fault Localization (SBFL) approaches are highly recognized due to their accuracy, efficiency, and simplicity. However, their effectiveness depends on the availability of fault-triggering tests. In the absence of such test failures, alternative approaches should be considered. In this paper, we investigate the use of stack traces for SBFL in the absence of fault-triggering tests. We study 60 crash reports from the Defects4J benchmark. We find that most of the studied crash reports do not contain fault-triggering tests. This results in very low efficiency when using traditional SBFL. On the other hand, when it comes to the stack traces, we find that in 98.3% of the studied bugs, the bugfix intention is directly correlated with the exception in the stack trace. In addition, in 78.3% of the bugs, the buggy method is directly reachable from the stack trace, with an average distance of only 0.34 method calls. This indicates that buggy methods are typically very close to the stack traces in terms of the execution call graph. Furthermore, our results demonstrate that even without employing advanced techniques, the stack traces alone provide a reliable indication of buggy locations, capable of locating more than half of the bugs in the Top 5. Finally, we develop a simplified SBFL method called SBEST that utilizes stack trace information as a proxy for failing tests, thereby integrating it with coverage information to perform SBFL.
SBEST located 17 bugs in Top-1, 32 in Top-3, and 33 in Top-5, with MAP and MRR values that are 32.22% and 17.43% higher than the Stack Trace ranking, respectively. The results suggest that combining coverage information with stack traces can enhance fault localization by offering a clearer view of system execution when fault-triggering tests are unavailable. Future studies can benefit from the findings of this research to develop more intricate techniques. In addition, all the data that we have made available can help studies in areas such as automatic bug fixing, test generation, and, of course, fault localization. As potential future work, we plan to expand our analysis by increasing our bug dataset by utilizing other sources beyond Defects4J.
8 Declarations
Conflicts of Interest.
All authors are from Concordia University and the University of Alberta. There may be conflicts with individuals from these two institutions.
Funding.
The authors have no funding to report.
Data Availability Statements.
We make the data publicly available online for replication and encourage future research in this area (Rafi et al., 2024).